OpenStudy (anonymous):

Solve this inqualite. |t + 4| > 10 Explain

OpenStudy (anonymous):

hello you need help

OpenStudy (anonymous):

Yes and fast plzzz I will medal and fan

OpenStudy (anonymous):

Okay! So you know what the absolute value is right? Or do you not?

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htm look this up it will help

OpenStudy (anonymous):

I need to know what the answer to the equations is and how did you get the answer but make it short

OpenStudy (anonymous):

Basically, t is anything from 7 or above. 6 + 4 = 11 which is greater than 10, and the absolute value is also 11, still greater.

OpenStudy (anonymous):

Solve (x – 3)(x – 4) = 0. Okay, this one is already factored for me. But how do I solve this? Think: If I multiply two things together and the result is zero, what can I say about those two things? I can say that at least one of them must also be zero. That is, the only way to multiply and get zero is to multiply by zero. (This is sometimes called "The Zero Factor Property" or "Rule" or "Principle".)

OpenStudy (anonymous):

Sorry, I meant 7 + 4 = 11 :P :)

OpenStudy (anonymous):

is it a b c or d

OpenStudy (anonymous):

8. Which expression is equivalent to |x| > 11?

OpenStudy (anonymous):

its quick and easy

OpenStudy (anonymous):

NO khan academy plzz

OpenStudy (anonymous):

why

OpenStudy (anonymous):

this is not kahn acadomy

OpenStudy (anonymous):

I fell asleep last time and all day I had a cramp

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you just help me on here

OpenStudy (anonymous):

sad

OpenStudy (anonymous):

did that help

OpenStudy (anonymous):

or no

OpenStudy (anonymous):

Dont be I just went on one of these and IT HELPED ALOT

OpenStudy (anonymous):

whats lat sopposed to mean to me sorry

OpenStudy (anonymous):

It helped and I am thanking you